OpenStudy (anonymous):

What are the solutions of 5x2 + 2x – 8 = 0?

OpenStudy (anonymous):

state clearly

OpenStudy (anonymous):

you need to use the quadric equation to solve

OpenStudy (anonymous):

quadratic formula *

OpenStudy (anonymous):

5x^2+2x-8....like this?

OpenStudy (zzr0ck3r):

yes like that. you do not have to use the quadratic formula, you can compleete the square like this x^2 + (2/5)x-(8/5) = 0 (x+(1/5))^2 = 8/5 + (2/5)^2 (x+(1/5)^2 = 8/5+ 4/25 (x+(1/5))^2 = 44/25 x+(1/5) = +-sqrt(44)/55 = 2sqrt(11)/5 x = 2sqrt(11)/5-(1/5) or x = -2sqrt(11)/5 - (1/5)

OpenStudy (anonymous):

ok